OpenStudy (anonymous):

using property of exponents, solve 32^t+1 = 16^1-t

OpenStudy (helder_edwin):

i guess it is \[\large 32^{t+1}=16^{1-t}. \] Replace 32=2^5 and 16=2^4. Then do some algebra (exponents' laws), u will get \[\large 2^A=2^B \] and from this u imply A=B.

OpenStudy (triciaal):

to continue 2^5(t+1) = 2^4(1-t) therefore 5(t + 1) = 4(1-t) 5t + 4t = 4 -5

OpenStudy (anonymous):

thank you, I got!

OpenStudy (triciaal):

you are welcome